OpenStudy (anonymous):

The Noodles & Company location in Parker, Colorado, sells its Pad Thai for $4.25 for a small, $5.25 for a regular with an additional $2.00 for protein for either size. If Noodles & Company is making a 300% markup based on the cost on their food, what is the corresponding markup based on selling price?

OpenStudy (cruffo):

300% x original cost = sell price 300% = 3.00 (as a decimal), so 3 x original cost = sell price If we divide both sides by 3 original cost = 1/3 x sell price 1/3 = 0.33333.... = about 33.3% (round to whatever you need) so the "corresponding markup based on selling price" is 33.3% ??? Not exactly a "markup", but I think that's what they are looking for.

OpenStudy (anonymous):

Thank you so much.